# Structural characterizations and Obstructions to Planar-Rips complexes

### v1

@@ 3.2. Perpendicular bisector
Figure 6. An illustration of the case when the ray $xP^{\rightarrow}$ intersects the side $cd$, and the intersection is outside the segment $\overline{xP}$. This yields a $3$-simplex $\left\{x,\tilde{P},c,d\right\}$ containing the point $P$.
Using Lemma 3.10, it suffices to show that for every vertex $P$ of $\mathcal{S}(\widetilde{\mathcal{O}})$, all the intersections of the ray $xP^{\rightarrow}$ with $\mathcal{S}(\widetilde{\mathcal{O}})$ lies between $x$ and $P$. On contrary, suppose there exists a vertex $P$ such that the extension of ray $xP^{\rightarrow}$ intersects the polygon $\mathcal{S}(\widetilde{\mathcal{O}})$, say at some edge $cd$ (See Figure 6). Then we have $P\in\mathrm{conv}\left\{x,c,d\right\}$. Now consider the antipodal point $\tilde{P}$ of $P$ in $\widetilde{\mathcal{O}}$ where the points $x,c,d$ are adjacent to $\tilde{P}$ i.e. $\left\{x,c,d\right\}\subset Lk(\tilde{P})$. Hence from Lemma 2.6, we have $P\in Lk(\tilde{P})$ i.e. the points $P$ and $\tilde{P}$ are adjacent which is a contradiction.
Now consider any $(n-1)$-dimensional simplex $\Delta$ in $\widetilde{\mathcal{O}}$ which contains the vertices $a$ and $b$. Since $\mathcal{R}(X)$ is an $n$-dimensional pseudomanifold, the $(n-1)$-dimensional simplex $\Delta$ is shared by exactly two $n$-dimensional simplices, one of which is $\Delta\cup\left\{x\right\}$. Hence, there exists another point $y$ outside $\widetilde{\mathcal{O}}$ such that the vertices of $\Delta$ are adjacent to $y$ i.e. $\Delta\cup\left\{y\right\}$ is an $n$-dimensional simplex in $\mathcal{R}(X)$. Note that $x$ and $y$ are non-adjacent since otherwise $\Delta\cup\left\{x,y\right\}$ forms an $(n+1)$-simplex.
@@
Figure 7. The points $x$ and $y$ are guard points of polygon $\tilde{\mathcal{O}}$. The adjacency of $y$ with $\tilde{a}$ $\tilde{b}$, and $y$ being a guard point of polygon $\tilde{\mathcal{O}}$ along the side $\overline{\tilde{a}\tilde{b}}$ prompts $y$ to be adjacent to every other vertex of $\tilde{\mathcal{O}}$. On the right, the segments $\overline{ya}$ and $\overline{z\tilde{a}}$ intersect, implying adjacency between $y$ and $z$.
Firstly, the point $y$ is adjacent to $a$ and $b$. Now consider the antipodal points $\tilde{a}$ and $\tilde{b}$ of $a$ and $b$ respectively. With the points $a,b,\tilde{a},\tilde{b},x,y$, we are in the setting of Lemma 3.14 with the parameter $r=1$, and the quadrilateral being $ab\tilde{a}\tilde{b}$. Hence from Lemma 3.14, the point $y$ is adjacent to $\tilde{a}$ and $\tilde{b}$, and is a guard point of quadrilateral $ab\tilde{a}\tilde{b}$ along the side $\tilde{a}\tilde{b}$. The induced structure is illustrated in Figure 7(a).
Now let $z$ be any point of $\widetilde{\mathcal{O}}$ other than $a,b,\tilde{a},\tilde{b}$. Since $y$ is a guard point of the quadrilateral $ab\tilde{a}\tilde{b}$, either the segment $\overline{ya}$ intersects the segment $\overline{z\tilde{a}}$ or the segment $\overline{yb}$ intersects the segment $\overline{z\tilde{b}}$ in $\mathcal{S}(X)$, depending on whether the point is on the left side or the right side of line $L_{xy}$ (For instance, see Figure 7(b)). For both the cases, from Lemma 2.4, either $y$ or $z$ is an apex since the antipodal points $a$ and $\tilde{a}$ or $b$ and $\tilde{b}$ cannot be adjacent i.e. $y$ is adjacent to $z$. ∎
Hence, the subcomplex spanned by $x$, $y$ and vertices of $\widetilde{\mathcal{O}}$ is isomorphic to $\mathcal{O}_{n}$. Since the only subcomplex of a pseudomanifold which is also a pseudomanifold of same dimension is itself, the Rips complex $\mathcal{R}(X)$ is isomorphic to $\mathcal{O}_{n}$..
@@
Let $K$ be an iterated chain of $\mathcal{O}_{n}$s, then $K$ is homotopy equivalent to wedge sum of circles and $n$-spheres i.e. $K\simeq\bigvee^{m}{\mathbb{S}^{n}}\vee\bigvee^{p}{\mathbb{S}^{1}}$ for some $p\geq 0$, where $m=\#$ of strongly connected components in $K$.
@@
Given $X\subset\mathbb{R}^{2}$, if the planar-Rips complex $\mathcal{R}(X)$ is an $n$-dimensional weak-pseudomanifold then $\mathcal{R}(X)$ is isomorphic to an iterated chain of $\mathcal{O}_{n}$s.
@@
Firstly, each strongly connected component is isomorphic to $\mathcal{O}_{n}$ since the corresponding induced subcomplex on each component inherits a planar-Rips structure; denote the strongly connected components by $\{P_{i}\}_{i=1}^{m}$. Now, suppose any two strongly connected components, say $P_{i}$ and $P_{j}$ have a non-empty intersection, then their intersection is a subcomplex of codimension at least $2$. Further, from Lemma 3.9, any $k$-simplex ($k\geq 2$) in $P_{i}$ and $P_{j}$ is non-degenerate, and the polygons $\mathcal{S}(P_{i})$ and $\mathcal{S}(P_{j})$ are convex; hence the intersection $P_{i}\cap P_{j}$ is a vertex when $n=2$, and is either a vertex or an edge when $n>2$.
@@
for some $p\geq 0$, where $m=\#$ of strongly connected components in $\mathcal{R}(X)$.
Note that $p=0$ unless a subcollection of strongly connected components can be arranged in a cyclic manner. Additionally, if $p\neq 0$ then $m>3$.
@@ 4. Characterization of minimal $2$-cycles and two-dimensional closed planar-Rips structures
In this section, we will prove C, laying the groundwork for recognizing two-dimensional, pure and closed planar-Rips structures, a more general class of $2$-pseudomanifolds. We will achieve this by first establishing its inherent local structure and showing that a minimal 2-cycle admitting a planar-Rips structure is isomorphic to $\mathcal{O}_{2}$.
@@
  • (a)
    A finite $n$-dimensional pure simplicial complex is said to be closed if every simplex of codimension $1$ has degree at least $2$ i.e. every $(n-1)$-simplex is contained in at least two $n$-simplices.
  • (b)
@@
For $X\subset\mathbb{R}^{2}$, if the planar-Rips complex $\mathcal{R}(X)$ is an $n$-dimensional pure and closed simplicial complex ($n\geq 2$), then every simplex is non-degenerate, meaning that no vertex in a simplex is contained within the convex hull of its remaining vertices.
@@
Assume the contrary i.e. let there be a degenerate simplex, say $\sigma$ of dimension at least $2$. Then there exists a vertex $v$ which is contained in the convex hull of remaining vertices of $\sigma$. Let $\tau$ be a maximal simplex containing $\sigma$ then $v\in\mathrm{conv}((\tau)-v)$. Since $\mathcal{R}(X)$ is closed, there exists another $n$-simplex $\gamma$ that does not contain $v$, such that $\gamma\neq\tau$ and $\tau-v\subset\gamma$. Now from Lemma 2.6, $v\cup\gamma$ is an $(n+1)$-simplex in $\mathcal{R}(X)$ which is not possible. ∎
@@
A specific arrangement of four points in $\mathcal{R}(X)$ as an induced subcomplex that contains exactly two $2$-simplices, such that the $1$-simplex shared by two $2$-simplices lies on one side of the line joining the two non-adjacent vertices.
Figure 8. $\boldsymbol{\gamma_{2}}$-configuration: The highlighted simplex (in black) denotes the base of the configuration.
Additionally, the $1$-simplex shared between two $2$-simplices in the $\gamma_{2}$-configuration will be referred to as the base of the configuration. The following result is a simple consequence of Lemma 4.2.
@@
[…]
@@
We will repeatedly use Proposition 4.4.$(ii)$ to prove this theorem. Firstly, as follows from this proposition, we start by taking a $\gamma_{2}$-configuration, say ${\left\langle{ABCD}\right\rangle}$ as an induced subcomplex in $\mathcal{R}(X)$ where the simplices $[ABC]$ and $[ABD]$ are in $\mathcal{R}(X)$, the points $C$ and $D$ are non-adjacent, and the segments $\overline{AD}$ and $\overline{BC}$ intersect. Note that the points $C$ and $D$ are on the same side of the line $L_{AB}$ i.e. without loss of generality, assume $C,D\in L^{+}_{AB}$.
@@
Figure 10. An illustration of 1. When $\mathcal{R}(X)$ is closed, $C$ and $D$ lie on same side of the bisector $L_{A|B}$, which is possile only when $C$ and $D$ are adjacent
@@
(a) $P\in L^{+}_{A|B}\cap\ cl(L^{+}_{BC})\cap L^{-}_{DC}$
(b) $P\in L^{+}_{A|B}\cap\ cl(L^{+}_{BC})\cap L^{+}_{DC}$
Figure 11. An illustration of proof of 2: For a $\gamma_{2}$-configuration ${\left\langle{ABCD}\right\rangle}$, the two regions are illustrated where point $P$ cannot lie.
We have $P\in L^{+}_{A|B}$. Now for any point $Z\in L^{+}_{A|B}\cap\ cl(L^{+}_{BC})$, there are two possibilities:
@@
In particular, if $P\in L^{+}_{A|B}\cap(L_{BC}\cup L^{+}_{BC})$ then either $C$ is adjacent to $D$ or $P$ is adjacent to $B$, neither of which is true. This is illustrated with an example in Figure 11(a) and Figure 11(b). Hence, $P\in L^{-}_{BC}$. With analogous arguments, $Q\in L^{+}_{AD}$. Since $P\in L^{+}_{A|B}\cap L^{-}_{BC}\subset L^{+}_{AD}$ and $Q\in L^{-}_{A|B}\cap L^{+}_{AD}\subset L^{-}_{BC}$, therefore $P,\ Q\in L^{+}_{AD}\cap L^{-}_{BC}$.
@@
For $X\subset\mathbb{R}^{2}$, if $\mathcal{R}(X)$ is a minimal $2$-cycle then it is isomorphic to $\mathcal{O}_{2}$.
@@
Theorem 3.15 for $n=2$ case is now simply a consequence of Theorem 4.7 since a pseudomanifold is a minimal cycle. In general, we like to know if Theorem 4.7 can be extended even for higher dimensions.
@@
The proof of Theorem 4.7 has a further implication that in a two-dimensional, pure and closed Rips complex, every $\gamma_{2}$-configuration further extends to $\mathcal{O}_{2}$, and is contained in precisely one copy of $\mathcal{O}_{2}$ in $\mathcal{R}(X)$. Hence, every $\gamma_{2}$-configuration is a part of precisely one non-trivial homology $2$-cycle. This further leads us to characterize all two-dimensional, pure and closed Rips complexes by considering the projection map $p:\mathcal{R}(X)\to\mathcal{S}(X)$.
@@
[…]
@@
Since $\mathcal{R}(X)$ contains a copy of $\mathcal{O}_{2}$ as an induced subcomplex, it suffices to show that there does not exist any other $2$-simplex outside or on the boundary edges of the induced subcomplex $\mathcal{O}_{2}$. But this follows from Proposition 4.4 and Lemma 4.11. ∎
@@
For the planar-Rips structure of boundary complex of octahedron $\mathcal{O}_{2}$, consider its shadow $\mathcal{S}(\mathcal{O}_{2})$ which is a convex hexagon under the projection map $p:\mathcal{R}(X)\to\mathcal{S}(X)$. Then $p$ is injective on the boundary of $\mathcal{O}_{2}$, and the inverse image of every point in the interior of hexagon $\mathcal{S}(\mathcal{O}_{2})$ has exactly two points in $\mathcal{O}_{2}$. More generally, it simply follows from Lemma 4.2 and Proposition 4.4.$ii$ that for a $2$-dimensional pure and closed Rips complex $\mathcal{R}(X)$, the inverse image of every point in the shadow $\mathcal{S}(X)$ has at most two points in $\mathcal{R}(X)$. Alternatively from an algorithmic perspective, each copy of $\mathcal{O}_{2}$ can be identified as the tripartite graph $K_{2,2,2}$ (See Section 5.1) in the graph of $1$-skeleton $\mathcal{R}^{(1)}(X)$.
@@
In a finite, two-dimensional, pure and closed planar-Rips complex $\mathcal{R}(X)$, denote $S$ to be set of all points in the shadow $\mathcal{S}(X)$ whose inverse image under the projection map $p:\mathcal{R}(X)\to\mathcal{S}(X)$ has exactly two points. Then
@@
    $b_{2}(\mathcal{R}(X))=b_{0}(S)$ i.e. the second Betti number $b_{2}(\mathcal{R}(X))$=# connected components in $S$. More specifically, $b_{0}$=Number of cycles of length $6$ in $\mathcal{R}(X)$ which has at least two of its diagonals intersecting at their interiors.
@@
Suppose $T$ is any connected component of $S$, then there exists two edges in $\mathcal{R}(X)$ whose projected interiors intersect in $T$. Hence from Proposition 4.4.$ii$, the two edges induce a $\gamma_{2}$-configuration, and from Theorem 4.7, this configuration further extends to an induced subcomplex of $\mathcal{O}_{2}$. Since the projected interior of the hexagon $\mathcal{S}(\mathcal{O}_{2})$ is connected, $\mathcal{S}(\mathcal{O}_{2})\subseteq T$. However, the map $p$ is injective on the boundary edges of $\mathcal{O}_{2}$ which implies $T=\mathcal{S}(\mathcal{O}_{2})$ i.e. every connected component of $S$ corresponds to an induced subcomplex of $\mathcal{O}_{2}$ in $\mathcal{R}(X)$. Hence, $b_{2}(\mathcal{R}(X))\geq b_{0}(S)$. Further, since any minimal $2$-cycle in $\mathcal{R}(X)$ is isomorphic to $\mathcal{O}_{2}$ (Corollary 4.8), it follows that $b_{2}(\mathcal{R}(X))=b_{0}(S)$. This proves (i). Now (ii) is a direct consequence of (i).
5. Obstructions to planar-Rips complexes and correlations with Unit Disk Graphs
In this section, we introduce the concept of obstructions to planar-Rips complexes, an area that remains unexplored in literature. However, it’s worth noting that the idea of ‘obstructions’ is not entirely novel [21, 22]. Armed with a comprehensive list of forbidden planar-Rips structures, we present a few examples that illustrate a potential framework for extracting obstructions to planar-Rips complex from these forbidden structures.
We begin with unifying the study of planar-Rips complexes and unit disk graphs by establishing a one-to-one correspondence between both the categories. The functors that facilitate this correspondence are “taking clique complex” and “taking $1$-skeleton”, which transform a unit disk graph to a planar-Rips complex and vice versa, respectively. This paves the way to retrieve the known obstructions to planar-Rips complexes in terms of their graph counterpart, called minimal non unit disk graphs. Additionally, we consolidate our findings on the classification of planar-Rips complexes from previous sections in terms of unit disk graphs, proving D.
Definition 5.1 (Unit disk graph).
A graph where each vertex represents a point in the Euclidean plane and an edge is drawn between two points if the distance between two points is less than a fixed threshold $r>0$.
Planar-Rips complexes are flag complexes that exhibit a key property: the clique complex of the graph of its $1$-skeleton is same as the original complex. It is easy to see that every unit disk graph can be identified as the $1$-skeleton of a planar-Rips complex. In fact, the following equivalence can be inferred.

### v2

@@ 3.2. Perpendicular bisector
Figure 6. An illustration of the case when the ray $\overrightarrow{xP}$ intersects the side cdcd, and the intersection is outside the segment $\overline{xP}$. This yields a $3$-simplex $\left\{x,\tilde{P},c,d\right\}$ containing the point $P$.
[…]
Now consider any (n1)(n-1)-dimensional simplex $\Delta$ in $\widetilde{\mathcal{O}}$ which contains the vertices $a$ and $b$. Since $\mathcal{R}(X)$ is an $n$-dimensional pseudomanifold, the (n1)(n-1)-dimensional simplex $\Delta$ is shared by exactly two $n$-dimensional simplices, one of which is $\Delta\cup\left\{x\right\}$. Hence, there exists another point $y$ outside $\widetilde{\mathcal{O}}$ such that the vertices of $\Delta$ are adjacent to $y$ i.e., $\Delta\cup\left\{y\right\}$ is an $n$-dimensional simplex in $\mathcal{R}(X)$. Note that $x$ and $y$ are non-adjacent since otherwise $\Delta\cup\left\{x,y\right\}$ forms an $(n+1)$-simplex.
@@
Figure 7. The points $x$ and $y$ are guard points of polygon $\tilde{\mathcal{O}}$. The adjacency of $y$ with $\tilde{a} &$ $\tilde{b}$, and $y$ being a guard point of polygon $\tilde{\mathcal{O}}$ along the side $\overline{\tilde{a}\tilde{b}}$ prompts $y$ to be adjacent to every other vertex of $\tilde{\mathcal{O}}$. On the right, the segments $\overline{ya}$ and $\overline{z\tilde{a}}$ intersect, implying adjacency between $y$ and $z$.
Firstly, the point $y$ is adjacent to $a$ and $b$. Now consider the antipodal points $\tilde{a}$ and $\tilde{b}$ of $a$ and $b$ respectively. With the points $a,b,\tilde{a},\tilde{b},x,y$, we are in the setting of Lemma 3.14 with the parameter $\delta=1$, and the quadrilateral being $ab\tilde{a}\tilde{b}$. Hence from Lemma 3.14, the point $y$ is adjacent to $\tilde{a}$ and $\tilde{b}$, and is a guard point of quadrilateral $ab\tilde{a}\tilde{b}$ along the side $\tilde{a}\tilde{b}$. The induced structure is illustrated in Figure 7(a).
Now let $z$ be any point of $\widetilde{\mathcal{O}}$ other than $a,b,\tilde{a},\tilde{b}$. Since $y$ is a guard point of the quadrilateral $ab\tilde{a}\tilde{b}$, either the segment $\overline{ya}$ intersects the segment $\overline{z\tilde{a}}$ or the segment $\overline{yb}$ intersects the segment $\overline{z\tilde{b}}$ in $\mathcal{S}(X)$, depending on whether the point is on the left side or the right side of line $L_{xy}$ (For instance, see Figure 7(b)). For both the cases, from Lemma 2.4, either $y$ or $z$ is an apex since the antipodal points $a$ and $\tilde{a}$ or $b$ and $\tilde{b}$ cannot be adjacent i.e., $y$ is adjacent to $z$. This concludes the proof of Claim 3. ∎
Hence, the subcomplex spanned by $x$, $y$ and vertices of $\widetilde{\mathcal{O}}$ is isomorphic to $\mathcal{O}_{n}$. Since the only subcomplex of a pseudomanifold which is also a pseudomanifold of same dimension is itself, the Rips complex $\mathcal{R}(X)$ is isomorphic to $\mathcal{O}_{n}$.
@@
Let $K$ be an iterated chain of $\mathcal{O}_{n}$s, then $K$ is homotopy equivalent to wedge sum of circles and $n$-spheres i.e., $K\simeq\bigvee^{m}{\mathbb{S}^{n}}\vee\bigvee^{p}{\mathbb{S}^{1}}$ for some $p\geq 0$, where m=#m=\# of strongly connected components in $K$.
@@
For any $n\geq 2$ and a finite subset $X\subset\mathbb{R}^{2}$, if the planar-Rips complex $\mathcal{R}(X)$ is an $n$-dimensional weak-pseudomanifold then $\mathcal{R}(X)$ is isomorphic to an iterated chain of $\mathcal{O}_{n}$s.
@@
Firstly, each strongly connected component is isomorphic to $\mathcal{O}_{n}$ since the corresponding induced subcomplex on each component inherits a planar-Rips structure; denote the strongly connected components by $\{P_{i}\}_{i=1}^{m}$ (See Figure 8 for an illustration). Now, suppose any two strongly connected components, say $P_{i}$ and $P_{j}$ have a non-empty intersection, then their intersection is a subcomplex of codimension at least $2$. Further, from Lemma 3.9, any $k$-simplex ($k\geq 2$) in $P_{i}$ and $P_{j}$ is non-degenerate, and the polygons $\mathcal{S}(P_{i})$ and $\mathcal{S}(P_{j})$ are convex; hence the intersection $P_{i}\cap P_{j}$ is a vertex when $n=2$, and is either a vertex or an edge when $n>2$.
@@
for some $p\geq 0$, where m=#m=\# of strongly connected components in $\mathcal{R}(X)$.
Remark 3.20.
Note that $p=0$ unless a subcollection of strongly connected components can be arranged in a “cyclic” manner, in which case $p=1$. Figure 8(b) illustrates how such a cyclic structure would look; however, it remains to be determined whether such a structure is possible. Additionally, if $p\neq 0$ then $m>3$.
(a) $m=4$, $p=0$
(b) $m=10$, $p=1$
Figure 8. Illustration of iterated chain of $\mathcal{O}_{n}$s for $n=2$. Correspondingly, the values of $m$ and $p$ are specified as in Corollary 3.19.
@@ 4. Characterization of minimal $2$-cycles and two-dimensional closed planar-Rips structures
In this section, we will prove Theorem C, laying the groundwork for recognizing two-dimensional, pure and closed planar-Rips structures, a more general class of $2$-pseudomanifolds. We will achieve this by first establishing its inherent local structure and showing that a minimal 2-cycle admitting a planar-Rips structure is isomorphic to $\mathcal{O}_{2}$.
@@
  1. (a)
    A finite $n$-dimensional pure simplicial complex is said to be closed if every simplex of codimension $1$ has degree at least $2$ i.e., every (n1)(n-1)-simplex is contained in at least two $n$-simplices.
  2. (b)
@@
For $X\subset\mathbb{R}^{2}$, if the planar-Rips complex $\mathcal{R}(X)$ is an $n$-dimensional pure and closed simplicial complex ($n\geq 2$), then every simplex is non-degenerate, meaning that no vertex in a simplex is contained within the convex hull of its remaining vertices in the shadow complex $\mathcal{S}(X)$.
@@
Assume the contrary i.e., let there be a degenerate simplex, say $\sigma$ of dimension at least $2$. Then there exists a vertex $v$ which is contained in the convex hull of remaining vertices of $\sigma$. Let $\tau$ be a maximal simplex containing $\sigma$ then $v\in\mathrm{conv}((\tau)-v)$. Since $\mathcal{R}(X)$ is closed, there exists another $n$-simplex $\gamma$ that does not contain $v$, such that $\gamma\neq\tau$ and $\tau-v\subset\gamma$. Now from Lemma 2.6, $v\cup\gamma$ is an $(n+1)$-simplex in $\mathcal{R}(X)$ which is not possible. ∎
@@
A $\boldsymbol{\gamma_{2}}$-configuration is an induced subcomplex of $\mathcal{R}(X)$ spanned by four vertices, which contains precisely two $2$-simplices, and the edge shared by them lies on one side of the line joining the two non-adjacent vertices in the shadow $\mathcal{S}(X)$.
Figure 9. An illustration of $\gamma_{2}$-configuration; The highlighted edge (in black) denotes the base of the configuration.
Additionally, the edge($1$-simplex) shared between two $2$-simplices in the $\gamma_{2}$-configuration will be referred to as the base of the configuration. The following result is a simple consequence of Lemma 4.2.
@@
Figure 10. Illustration of all possible and forbidden cases in the proof of Proposition 4.4.(ii).
@@
We will repeatedly use Proposition 4.4.(ii) to prove this theorem. Firstly, as follows from this proposition, we start by taking a $\gamma_{2}$-configuration, say $\left\langle{ABCD}\right\rangle$} as an induced subcomplex in $\mathcal{R}(X)$ where the simplices $[ABC]$ and $[ABD]$ are in $\mathcal{R}(X)$, the points $C$ and $D$ are non-adjacent, and the segments $\overline{AD}$ and $\overline{BC}$ intersect. Note that the points $C$ and $D$ are on the same side of the line $L_{AB}$ i.e., without loss of generality, assume $C,D\in L^{+}_{AB}$.
@@
Figure 11. An illustration of Claim 1: Here, the points $C$ and $D$ are taken to be on the left side of the bisector $L_{A|B}$. Corresponding to this arrangement, $Q$ lies in the hatched region, implying that $D\in\mathrm{conv}\left\{B,C,Q\right\}$, as represented by the solid shaded region.
@@
(a) $P\in L^{+}_{A|B}\cap\ \mathrm{cl}(L^{+}_{BC})\cap L^{-}_{DC}$
(b) $P\in L^{+}_{A|B}\cap\ \mathrm{cl}(L^{+}_{BC})\cap L^{+}_{DC}$
[…]
@@
[…]
@@
For a finite subset $X\subset\mathbb{R}^{2}$, if $\mathcal{R}(X)$ is a minimal $2$-cycle then it is isomorphic to $\mathcal{O}_{2}$.
@@
It can be seen that Theorem 3.15 for $n=2$ case is now simply a consequence of Theorem 4.7 since a pseudomanifold is a minimal cycle. In general, we like to know if Theorem 4.7 can be extended even for higher dimensions.
@@
The proof of Theorem 4.7 has a further implication that in a two-dimensional, pure and closed Rips complex, every $\gamma_{2}$-configuration further extends to $\mathcal{O}_{2}$, and is contained in precisely one copy of $\mathcal{O}_{2}$ in $\mathcal{R}(X)$. Hence, every $\gamma_{2}$-configuration is a part of precisely one non-trivial homology $2$-cycle. This further leads us to characterize all two-dimensional, pure and closed Rips complexes by considering the projection map $p\colon\mathcal{R}(X)\to\mathcal{S}(X)$.
@@
Suppose not, then there exists two boundary edges in $\mathcal{R}(X)$, say $[AB]$ and $[CD]$ whose projected interiors intersect in the shadow $\mathcal{S}(X)$ with $B$ and $C$ being non-adjacent. Now it follows from Proposition 4.4.(i) that the induced subcomplex $\left\langle{ABCD}\right\rangle$} is a $\gamma_{2}$-configuration, which further extends to $\mathcal{O}_{2}$ as shown in Theorem 4.7. But this is not possible since $[AB]$ and $[CD]$ are boundary edges in $\mathcal{R}(X)$.
@@
Since $\mathcal{R}(X)$ contains a copy of $\mathcal{O}_{2}$ as an induced subcomplex, it suffices to show that there does not exist any other 22-simplex outside or on the boundary edges of the induced subcomplex $\mathcal{O}_{2}$. But this follows from Proposition 4.4 and Lemma 4.11. ∎
@@
For the planar-Rips structure of boundary complex of octahedron $\mathcal{O}_{2}$, consider its shadow $\mathcal{S}(\mathcal{O}_{2})$ which is a convex hexagon under the projection map $p\colon\mathcal{R}(X)\to\mathcal{S}(X)$. Then $p$ is injective on the boundary of $\mathcal{O}_{2}$, and the inverse image of every point in the interior of hexagon $\mathcal{S}(\mathcal{O}_{2})$ has exactly two points in $\mathcal{O}_{2}$. More generally, it simply follows from Lemma 4.2 and Proposition 4.4.(ii) that for a $2$-dimensional pure and closed Rips complex $\mathcal{R}(X)$, the inverse image of every point in the shadow $\mathcal{S}(X)$ has at most two points in $\mathcal{R}(X)$. Alternatively from an algorithmic perspective, each copy of $\mathcal{O}_{2}$ can be identified as the tripartite graph $K_{2,2,2}$ (See Section 5.1) in the graph of $1$-skeleton $\mathcal{R}^{(1)}(X)$.
@@
In a finite, two-dimensional, pure and closed planar-Rips complex $\mathcal{R}(X)$, denote $S$ to be set of all points in the shadow $\mathcal{S}(X)$ whose inverse image under the projection map $p\colon\mathcal{R}(X)\to\mathcal{S}(X)$ has exactly two points. Then
@@
    $b_{2}(\mathcal{R}(X))=b_{0}(S)$ i.e., the second Betti number $b_{2}(\mathcal{R}(X))$=# connected components in $S$. More specifically, $b_{0}$=Number of cycles of length $6$ in $\mathcal{R}(X)$ which has at least two of its diagonals intersecting at their interiors.
@@
Suppose $T$ is any connected component of $S$, then there exists two edges in $\mathcal{R}(X)$ whose projected interiors intersect in $T$. Hence from Proposition 4.4.(ii), the two edges induce a $\gamma_{2}$-configuration, and from Theorem 4.7, this configuration further extends to an induced subcomplex of $\mathcal{O}_{2}$. Since the projected interior of the hexagon $\mathcal{S}(\mathcal{O}_{2})$ is connected, $\mathcal{S}(\mathcal{O}_{2})\subseteq T$. However, the map $p$ is injective on the boundary edges of $\mathcal{O}_{2}$ which implies $T=\mathcal{S}(\mathcal{O}_{2})$ i.e., every connected component of $S$ corresponds to an induced subcomplex of $\mathcal{O}_{2}$ in $\mathcal{R}(X)$. Hence, $b_{2}(\mathcal{R}(X))\geq b_{0}(S)$. Further, since any minimal $2$-cycle in $\mathcal{R}(X)$ is isomorphic to $\mathcal{O}_{2}$ (Corollary 4.8), it follows that $b_{2}(\mathcal{R}(X))=b_{0}(S)$. This proves (i). Now (ii) is a direct consequence of (i). ∎
5. Correlations with unit disk graphs
In this section, we introduce the concept of obstructions of planar-Rips complexes, an area that remains unexplored in literature. However, it’s worth noting that the idea of ‘obstructions’ is not entirely novel [25, 28]. Armed with a wide spectrum of forbidden planar-Rips structures, we present a few examples that illustrate a potential framework for extracting obstructions of planar-Rips complex from these forbidden structures.
We begin with unifying the study of planar-Rips complexes and unit disk graphs by establishing a one-to-one correspondence between both the categories. The functors that facilitate this correspondence are “taking clique complex” and “taking $1$-skeleton”, which transform a unit disk graph to a planar-Rips complex and vice versa, respectively. For simplicity, we will assume all the graphs to be connected.
Definition 5.1 (Disk graph).
A disk graph is an intersection graph of equal radius disks in the Euclidean plane. Equivalently, a disk graph is a geometric graph where each vertex corresponds to a point in the Euclidean plane and an edge is drawn between two points if and only if the distance between them is less than a fixed threshold, say $\delta>0$.
Note that in case of $\delta=1$, they are commonly referred to as unit disk graphs. Planar-Rips complexes are flag complexes that exhibit a key property: the clique complex of the graph of its $1$-skeleton is same as the original complex. It is easy to see that every disk graph can be identified as the $1$-skeleton of a planar-Rips complex. In fact, the following equivalence can be inferred.